OpenStudy (anonymous):

f(x)= 2^e^sinx find derivative

hartnn (hartnn):

@Walker21 Hi :) \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) do you know the chain rule ?? also, first you can put 2^e = a (another constant) here....

OpenStudy (anonymous):

Did Goku just use instant transmission?

OpenStudy (anonymous):

lol nvm

OpenStudy (anonymous):

so chain rile first?

hartnn (hartnn):

you'll need chain rule for a^(sin x) because standard function is a^x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still lost

OpenStudy (anonymous):

Take e^sinx=u d(2^u)/dx= 2^u log2 du/dx Sloving for du/dx = d(e^sinx)/dx = e^sinx (cosx) So final answer is 2^(e^sinx).log2.e^sinx .cosx

OpenStudy (anonymous):

I hope it was clear ...

OpenStudy (anonymous):

thanks helps

OpenStudy (anonymous):

Medal is its accurate :)

OpenStudy (anonymous):

are you good at finding points on a graph at which the tagent line is horizonatal?

OpenStudy (anonymous):

Not so much. No

OpenStudy (anonymous):

ok